# Zeroth-Order Constrained Optimization from a Control Perspective via Feedback Linearization

### v1

@@ Appendix C Proof of Theorem 2
$\quad y_{t+1}-y_{t}=J_{h}(x_{t})(x_{t+1}-x_{t})+\epsilon_{t}$
$=-\eta_{t}J_{h}(x_{t})\left(\widetilde{\nabla}f(x_{t})+\widetilde{J}_{h}(x_{t})^{\top}\lambda_{t}\right)+\epsilon_{t}$
$=-\eta_{t}J_{h}(x_{t})\left(\widetilde{\nabla}f(x_{t})+\widetilde{J}_{h}(x_{t})^{\top}\lambda_{t}^{\star}\right)+\eta_{t}\underbrace{J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top}(\lambda_{t}^{\star}-\lambda_{t})}_{:=\Delta_{t}}+\epsilon_{t}$
$=-\eta_{t}(Kh(x_{t})+s^{\star})+\eta_{t}\Delta_{t}+\epsilon_{t}$
$=-\eta_{t}Ky_{t}-\eta^{t}s^{\star}+\eta_{t}\Delta_{t}+\epsilon_{t}$
@@
$\|[y_{t+1}]_{+}\|\leq(1-\eta_{t}{\lambda_{\min}}(K))\|[y_{t}]_{+}\|+\eta_{t}\|\Delta_{t}\|+C_{1}\eta_{t}^{2}$
Further, from Lemma 4, we have that
$\|\Delta_{t}\|\leq C_{2}r_{2}^{2},~~\textup{where }C_{2}=n^{2}\overline{L}_{h}^{2}R\left(\frac{4096n\overline{L}_{h}(L_{f}\overline{L}_{h}+\|K\|H)}{\underline{L}_{h}^{4}}+\frac{64L_{f}}{\underline{L}_{h}^{2}}\right)$
Thus the rest of the proof follows exactly the same derivation as the proof of Theorem 1, here we repeat as:
$\|[y_{t+1}]_{+}\|$ $\leq(1-\eta_{t}{\lambda_{\min}}(K))\|[y_{t}]_{+}\|+\eta_{t}C_{2}r_{2}^{2}+C_{1}\eta_{t}^{2}$
$\Longrightarrow\|[y_{t}]_{+}\|$ $\leq\prod_{s=0}^{t-1}(1-\eta_{s}{\lambda_{\min}}(K))\|[y_{0}]_{+}\|+C_{2}r_{2}^{2}\sum_{s=0}^{t-1}\prod_{\tau=s+1}^{t-1}(1-\eta_{\tau}{\lambda_{\min}}(K))\eta_{s}+C_{1}\sum_{s=0}^{t-1}\prod_{\tau=s+1}^{t-1}(1-\eta_{\tau}{\lambda_{\min}}(K))\eta_{s}^{2}$
@@
$\|[y_{t}]_{+}\|$ $\leq(1-\eta{\lambda_{\min}}(K))^{t}\|[y_{0}]_{+}\|+C_{2}r_{2}^{2}\sum_{s=0}^{t-1}(1-\eta{\lambda_{\min}}(K))^{s}\eta+C_{1}\eta^{2}\sum_{s=0}^{t-1}(1-\eta{\lambda_{\min}}(K))^{2}$
$\leq(1-\eta{\lambda_{\min}}(K))^{t}\|[y_{0}]_{+}\|+\frac{C_{2}r_{2}^{2}}{{\lambda_{\min}}(K)}+\frac{C_{1}\eta}{{\lambda_{\min}}(K)}$
@@
$\|[y_{t}]_{+}\|$ $\leq\prod_{s=0}^{t-1}\left(1-\frac{\eta{\lambda_{\min}}(K)}{\sqrt{s+1}}\right)\|[y_{0}]_{+}\|+C_{2}r_{2}^{2}\sum_{s=0}^{t-1}\prod_{\tau=s+1}^{t-1}\left(1-\frac{\eta{\lambda_{\min}}(K)}{\sqrt{\tau+1}}\right)\frac{\eta}{\sqrt{s+1}}$
$\qquad+C_{1}\sum_{s=0}^{t-1}\prod_{\tau=s+1}^{t-1}\left(1-\frac{\eta{\lambda_{\min}}(K)}{\sqrt{\tau+1}}\right)\frac{\eta^{2}}{s}$
@@
$T_{B}\geq 32\left(m\log\left(\frac{192\cdot n\cdot\overline{L}_{h}^{2}}{\underline{L}_{h}^{2}}\right)+\log\left(\frac{T_{G}}{\delta}\right)\right)\sim O\left(m\left(\log(n)+\log\left(\frac{\overline{L}_{h}}{\underline{L}_{h}}\right)\right)+\log\left(\frac{T_{G}}{\delta}\right)\right)$
and
$r_{1}\leq\frac{L_{h}}{8\sqrt{2\overline{L}_{h}R}},\quad r_{2}\leq\frac{L_{h}}{8\sqrt{2n\overline{L}_{h}R}}$
then with probability at least $1-\delta$,
$\|x_{t+1}-x_{t}\|\leq\eta_{t}\left(nL_{f}+\frac{64n\overline{L}_{h}(L_{f}\overline{L}_{h}+\|K\|H)}{\underline{L}_{h}^{2}}\right)$
@@
$\Longrightarrow$~~ $\|\widetilde{\nabla}f(x_{t})\|\leq nL_{f}.$
@@
$\|\widetilde{J}(x_{t})\|\leq n\overline{L}_{h},~~\|G_{f}\|\leq L_{f}\overline{L}_{h}.$
Further, from Lemma 5, when
$r_{1}\leq\frac{L_{h}}{8\sqrt{2\overline{L}_{h}R}},\quad r_{2}\leq\frac{L_{h}}{8\sqrt{2n\overline{L}_{h}R}},$
we have that
$\sigma_{\min}(G_{h})\geq\frac{\underline{L}_{h}^{2}}{64}.$
@@
$G_{h}\lambda_{t}+G_{f}=Kh(x_{t})+s,~~\lambda_{t}\geq 0,~~s\geq 0,~~\lambda_{t}^{\top}s=0.$
@@
$[\lambda_{t}]_{\mathcal{I}^{c}}$ $=0$
$[G_{h}]_{\mathcal{I}\mathcal{I}}[\lambda_{t}]_{\mathcal{I}}+[G_{h}-Kh(x_{t})]_{\mathcal{I}}$ $=0.$
@@
$\|\lambda_{t}\|=\left\|[G_{h}]^{-1}_{\mathcal{I}\mathcal{I}}[G_{f}-Kh(x_{t})]_{\mathcal{I}}\right\|$
From Cauchy’s Interlacing Theorem we get that
$\sigma_{\min}([G_{h}]_{\mathcal{I}\mathcal{I}})\geq\frac{\underline{L}_{h}^{2}}{64}$
Thus
$\|\lambda_{t}\|\leq\frac{64}{\underline{L}_{h}^{2}}\|G_{h}-Kh(x_{t})\|\leq\frac{64n\overline{L}_{h}(L_{f}\overline{L}_{h}+\|K\|H)}{\underline{L}_{h}^{2}}$
Finally
$\|x_{t+1}-x_{t}\|$ $\leq\eta_{t}(\|\widetilde{\nabla}f(x)\|+\|\widetilde{J}_{h}(x_{t})\|\|\lambda_{t}\|)$
$\leq\eta_{t}\left(nL_{f}+\frac{64n\overline{L}_{h}(L_{f}\overline{L}_{h}+\|K\|H)}{\underline{L}_{h}^{2}}\right),$
which completes the proof. ∎
Lemma 4.
We define the auxiliary variable $\lambda_{t}^{\star}$ as in equation 20. Under the conditions as stated in Lemma 3, we have that
$\|J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top}(\lambda_{t}^{\star}-\lambda_{t})\|\leq n^{2}\overline{L}_{h}^{2}R\left(\frac{4096nL_{f}\overline{L}_{h}^{2}}{\underline{L}_{h}^{4}}+\frac{64L_{f}}{\underline{L}_{h}^{2}}\right)r_{2}^{2}$
Proof.
Define
$~~~A=J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top},$ $\hskip-35.0ptb$ $=J_{h}(x_{t})\widetilde{\nabla}f(x_{t})-Kh(x)$
$\Delta A=G_{h}-J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top},$ $\hskip-35.0pt\Delta b$ $=G_{f}-J_{h}(x_{t})\widetilde{\nabla}f(x_{t})$
By Taylor’s expansion,
$\|\Delta A\|$ $=\|G_{h}-J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top}\|\leq n\overline{L}_{h}Rr_{2}^{2},$
$\|\Delta b\|$ $=\|G_{f}-J_{h}(x_{t})\widetilde{\nabla}f(x_{t})\|\leq nL_{f}Rr_{2}^{2},$
$\|G_{f}\|,$ $\|J_{h}(x_{t})\widetilde{\nabla}f(x_{t})\|\leq nL_{f}\overline{L}_{h}.$
$A(\alpha):=A+\alpha\Delta A,~~B(\alpha)=B+\alpha\Delta B$
and define $\lambda(\alpha)$ to be the solution of
$A(\alpha)\lambda(\alpha)+b(\alpha)=s(\alpha),~~\lambda(\alpha)\geq 0,~~s(\alpha)\geq 0,~~\lambda(\alpha)^{\top}s(\alpha)=0$ (21)
We can find a sequence of $\{\alpha_{i}\}$ such that $0=\alpha_{0}<\alpha_{1}<...<\alpha_{N}=1$ such that with in each interval $\alpha,\alpha^{\prime}\in(\alpha_{i},\alpha_{i+1})$, $\lambda(\alpha)$ and $\lambda(\alpha^{\prime})$ shares exactly the same support, which we denote as $\mathcal{I}_{i}$. And from equation 21 we know that for any $\alpha\in[\alpha_{1},\alpha_{2}]$, $\lambda(\alpha)$ can be written as follows:
$[\lambda(\alpha)]_{\mathcal{I}_{i}}=[A(\alpha)]_{\mathcal{I}_{i}\mathcal{I}_{i}}^{-1}b(\alpha),~~[\lambda(\alpha)]_{\mathcal{I}_{i}^{c}}=0$
And thus we get
$\quad\|\lambda(\alpha_{i+1})-\lambda(\alpha_{i})\|\leq\|[A(\alpha_{i+1})]_{\mathcal{I}_{i}\mathcal{I}_{i}}^{-1}b(\alpha_{i+1})-[A(\alpha_{i})]_{\mathcal{I}_{i}\mathcal{I}_{i}}^{-1}b(\alpha_{i})\|$
$\leq\|A(\alpha_{i+1})]_{\mathcal{I}_{i}\mathcal{I}_{i}}^{-1}\|\|A(\alpha_{i})]_{\mathcal{I}_{i}\mathcal{I}_{i}}^{-1}\|A(\alpha_{i+1})-A(\alpha_{i})\|\|\|b(\alpha_{i+1})\|+\|A(\alpha_{i})]_{\mathcal{I}_{i}\mathcal{I}_{i}}^{-1}\|\|b(\alpha_{i+1})-b(\alpha_{i})\|$
$=(\alpha_{i+1}-\alpha_{i})\left(\|A(\alpha_{i+1})]_{\mathcal{I}_{i}\mathcal{I}_{i}}^{-1}\|\|A(\alpha_{i})]_{\mathcal{I}_{i}\mathcal{I}_{i}}^{-1}\|\Delta A\|\|\|b(\alpha_{i+1})\|+\|A(\alpha_{i})]_{\mathcal{I}_{i}\mathcal{I}_{i}}^{-1}\|\|\Delta b\|\right)$
$\leq(\alpha_{i+1}-\alpha_{i})\left(\overline{L_{h}}\|A(\alpha_{i+1})]_{\mathcal{I}_{i}\mathcal{I}_{i}}^{-1}\|\|A(\alpha_{i})]_{\mathcal{I}_{i}\mathcal{I}_{i}}^{-1}\|\|b(\alpha_{i+1})\|+L_{f}\|A(\alpha_{i})]_{\mathcal{I}_{i}\mathcal{I}_{i}}^{-1}\|\right)nRr_{2}^{2}$
Further, from Lemma 5 and Cauchy’s interlacing theorem (cf. Hwang [2004]) we know that for any principal minor of $A(\alpha)$ we have
$\|[A(\alpha)^{-1}]_{\mathcal{I}\mathcal{I}}\|\leq\frac{64}{\underline{L}_{h}^{2}},~~\forall\alpha\in[0,1]$
and clearly
$\|b(\alpha)\|\leq nL_{f}\overline{L}_{h}+\|K\|H$
And thus we get
$\|\lambda(\alpha_{i+1})-\lambda(\alpha_{i})\|\leq(\alpha_{i+1}-\alpha_{i})\left(\frac{4096n\overline{L}_{h}(L_{f}\overline{L}_{h}+\|K\|H)}{\underline{L}_{h}^{4}}+\frac{64L_{f}}{\underline{L}_{h}^{2}}\right)nRr_{2}^{2}$

### v2

@@ Appendix A Other Related Works on Derivative Free Constrained Optimization
$\textstyle\quad y_{t+1}-y_{t}=J_{h}(x_{t})(x_{t+1}-x_{t})+\epsilon_{t}=-\eta_{t}J_{h}(x_{t})\left(\widetilde{\nabla}f(x_{t})+\widetilde{J}_{h}(x_{t})^{\top}\lambda_{t}\right)+\epsilon_{t}$
$\textstyle=-\eta_{t}J_{h}(x_{t})\left(\widetilde{\nabla}f(x_{t})+\widetilde{J}_{h}(x_{t})^{\top}\lambda_{t}^{\star}\right)+\eta_{t}\underbrace{J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top}(\lambda_{t}^{\star}-\lambda_{t})}_{:=\Delta_{t}}+\epsilon_{t}$
$\textstyle=-\eta_{t}(Kh(x_{t})+s^{\star})+\eta_{t}\Delta_{t}+\epsilon_{t}=-\eta_{t}Ky_{t}-\eta^{t}s^{\star}+\eta_{t}\Delta_{t}+\epsilon_{t}$
@@
$\textstyle\|[y_{t+1}]_{+}\|\leq(1-\eta_{t}{\lambda_{\min}}(K))\|[y_{t}]_{+}\|+\eta_{t}\|\Delta_{t}\|+C_{1}\eta_{t}^{2}$
Further, from Lemma D.4, we have that
$\textstyle\|\Delta_{t}\|\leq C_{2}r_{2}^{2},~~\textup{where }C_{2}=n^{2}\overline{L}_{h}^{2}R\left(\frac{4096n\overline{L}_{h}(L_{f}\overline{L}_{h}+\|K\|H)}{\underline{L}_{h}^{4}}+\frac{64L_{f}}{\underline{L}_{h}^{2}}\right)$
Thus the rest of the proof follows exactly the same derivation as the proof of Theorem 3.1, here we repeat as:
$\textstyle\|[y_{t+1}]_{+}\|$ $\textstyle\leq(1-\eta_{t}{\lambda_{\min}}(K))\|[y_{t}]_{+}\|+\eta_{t}C_{2}r_{2}^{2}+C_{1}\eta_{t}^{2}$
$\textstyle\Longrightarrow\|[y_{t}]_{+}\|$ $\textstyle\leq\prod_{s=0}^{t-1}(1-\eta_{s}{\lambda_{\min}}(K))\|[y_{0}]_{+}\|+C_{2}r_{2}^{2}\sum_{s=0}^{t-1}\prod_{\tau=s+1}^{t-1}(1-\eta_{\tau}{\lambda_{\min}}(K))\eta_{s}$
$\textstyle\qquad\qquad\qquad\qquad\qquad\qquad\qquad+C_{1}\sum_{s=0}^{t-1}\prod_{\tau=s+1}^{t-1}(1-\eta_{\tau}{\lambda_{\min}}(K))\eta_{s}^{2}$
@@
$\textstyle\|[y_{t}]_{+}\|$ $\textstyle\leq(1-\eta{\lambda_{\min}}(K))^{t}\|[y_{0}]_{+}\|+C_{2}r_{2}^{2}\sum_{s=0}^{t-1}(1-\eta{\lambda_{\min}}(K))^{s}\eta$
$\textstyle\quad+C_{1}\eta^{2}\sum_{s=0}^{t-1}(1-\eta{\lambda_{\min}}(K))^{2}$
$\textstyle\leq(1-\eta{\lambda_{\min}}(K))^{t}\|[y_{0}]_{+}\|+\frac{C_{2}r_{2}^{2}}{{\lambda_{\min}}(K)}+\frac{C_{1}\eta}{{\lambda_{\min}}(K)}$
@@
$\textstyle\|[y_{t}]_{+}\|$ $\textstyle\leq\prod_{s=0}^{t-1}\left(1-\frac{\eta{\lambda_{\min}}(K)}{\sqrt{s+1}}\right)\|[y_{0}]_{+}\|+C_{2}r_{2}^{2}\sum_{s=0}^{t-1}\prod_{\tau=s+1}^{t-1}\left(1-\frac{\eta{\lambda_{\min}}(K)}{\sqrt{\tau+1}}\right)\frac{\eta}{\sqrt{s+1}}$
$\textstyle\qquad+C_{1}\sum_{s=0}^{t-1}\prod_{\tau=s+1}^{t-1}\left(1-\frac{\eta{\lambda_{\min}}(K)}{\sqrt{\tau+1}}\right)\frac{\eta^{2}}{s}$
@@
$\textstyle\quad+C_{1}\eta^{2}e^{-\eta{\lambda_{\min}}(K)\sqrt{t}}\sum_{s=1}^{t}e^{\eta{\lambda_{\min}}(K)\sqrt{s}}\frac{1}{s}$
@@
[…]
and $r_{1}\leq\frac{L_{h}}{8\sqrt{2\overline{L}_{h}R}},\quad r_{2}\leq\frac{L_{h}}{8\sqrt{2n\overline{L}_{h}R}},$ then with probability at least $1-\delta$,
$\textstyle\|x_{t+1}-x_{t}\|\leq\eta_{t}\left(nL_{f}+\frac{64n\overline{L}_{h}(L_{f}\overline{L}_{h}+\|K\|H)}{\underline{L}_{h}^{2}}\right)$
@@
$\textstyle|f(x_{t}+r_{1}u_{i})-f(x_{t}-r_{1}u_{i})|\leq 2r_{1}\nabla f(x_{t}+\tilde{r}_{1}u_{i})^{\top}u_{i}\leq 2r_{1}L_{f},$
$\textstyle\Longrightarrow$ $\textstyle\|\widetilde{\nabla}f(x_{t})\|\leq nL_{f}.$
@@
$\textstyle\|\widetilde{J}(x_{t})\|\leq n\overline{L}_{h},~~\|G_{f}\|\leq L_{f}\overline{L}_{h}.$
Further, from Lemma E.1, when $r_{1}\leq\frac{L_{h}}{8\sqrt{2\overline{L}_{h}R}},\quad r_{2}\leq\frac{L_{h}}{8\sqrt{2n\overline{L}_{h}R}},$ we have that
$\textstyle\sigma_{\min}(G_{h})\geq\frac{\underline{L}_{h}^{2}}{64}.$
@@
$\textstyle G_{h}\lambda_{t}+G_{f}=Kh(x_{t})+s,~~\lambda_{t}\geq 0,~~s\geq 0,~~\lambda_{t}^{\top}s=0.$
@@
${}_{\mathcal{I}^{c}}=0,\quad[G_{h}]_{\mathcal{I}\mathcal{I}}[\lambda_{t}]_{\mathcal{I}}+[G_{h}-Kh(x_{t})]_{\mathcal{I}}=0$
$\textstyle\Longrightarrow~\|\lambda_{t}\|=\left\|[G_{h}]^{-1}_{\mathcal{I}\mathcal{I}}[G_{f}-Kh(x_{t})]_{\mathcal{I}}\right\|$
From Cauchy’s Interlacing Theorem we get that $\sigma_{\min}([G_{h}]_{\mathcal{I}\mathcal{I}})\geq\frac{\underline{L}_{h}^{2}}{64}$. Thus
$\textstyle\|\lambda_{t}\|\leq\frac{64}{\underline{L}_{h}^{2}}\|G_{h}-Kh(x_{t})\|\leq\frac{64n\overline{L}_{h}(L_{f}\overline{L}_{h}+\|K\|H)}{\underline{L}_{h}^{2}}$
$\textstyle\textup{Finally}~\|x_{t+1}\!-\!x_{t}\|\leq\eta_{t}(\|\widetilde{\nabla}f(x)\|\!+\!\|\widetilde{J}_{h}(x_{t})\|\|\lambda_{t}\|)\leq\eta_{t}\left(nL_{f}+\frac{64n\overline{L}_{h}(L_{f}\overline{L}_{h}+\|K\|H)}{\underline{L}_{h}^{2}}\right),$
which completes the proof.
Lemma D.4.
We define the auxiliary variable $\lambda_{t}^{\star}$ as in (20). Under the conditions as stated in Lemma D.2, we have that
$\textstyle\|J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top}(\lambda_{t}^{\star}-\lambda_{t})\|\leq n^{2}\overline{L}_{h}^{2}R\left(\frac{4096nL_{f}\overline{L}_{h}^{2}}{\underline{L}_{h}^{4}}+\frac{64L_{f}}{\underline{L}_{h}^{2}}\right)r_{2}^{2}$
Proof D.5.
Define
$\textstyle~~~A=J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top},$ $\textstyle b$ $\textstyle=J_{h}(x_{t})\widetilde{\nabla}f(x_{t})-Kh(x)$
$\textstyle\Delta A=G_{h}-J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top},$ $\textstyle\hskip-25.0pt\Delta b$ $\textstyle=G_{f}-J_{h}(x_{t})\widetilde{\nabla}f(x_{t})$
From Lemma F.1: $\|\Delta A\|\!\leq\!n\overline{L}_{h}Rr_{2}^{2},~\|\Delta b\|\!\leq\!nL_{f}Rr_{2}^{2},~\|G_{f}\|,\!\|J_{h}(x_{t})\widetilde{\nabla}f(x_{t})\|\!\leq\!nL_{f}\overline{L}_{h}.$
$\textstyle A(\alpha)=A+\alpha\Delta A,~~B(\alpha)=B+\alpha\Delta B$
and define $\lambda(\alpha)$ to be the solution of
(21) $\textstyle A(\alpha)\lambda(\alpha)+b(\alpha)=s(\alpha),~~\lambda(\alpha)\geq 0,~~s(\alpha)\geq 0,~~\lambda(\alpha)^{\top}s(\alpha)=0$
We can find a sequence of $\{\alpha_{i}\}$ such that $0=\alpha_{0}<\alpha_{1}<...<\alpha_{N}=1$ such that with in each interval $\alpha,\alpha^{\prime}\in(\alpha_{i},\alpha_{i+1})$, $\lambda(\alpha)$ and $\lambda(\alpha^{\prime})$ shares exactly the same support, which we denote as $\mathcal{I}_{i}$. And from (21) we know that for any $\alpha\in[\alpha_{1},\alpha_{2}]$, $\lambda(\alpha)$ can be written as follows:
${}_{\mathcal{I}_{i}}=[A(\alpha)]_{\mathcal{I}_{i}\mathcal{I}_{i}}^{-1}b(\alpha),~~[\lambda(\alpha)]_{\mathcal{I}_{i}^{c}}=0$
And thus we get
$\textstyle\quad\|\lambda(\alpha_{i+1})-\lambda(\alpha_{i})\|\leq\|[A(\alpha_{i+1})]_{\mathcal{I}_{i}\mathcal{I}_{i}}^{-1}b(\alpha_{i+1})-[A(\alpha_{i})]_{\mathcal{I}_{i}\mathcal{I}_{i}}^{-1}b(\alpha_{i})\|$
$\textstyle\leq\!\|\!A(\!\alpha_{i+1}\!)]_{\mathcal{I}_{i}\mathcal{I}_{i}}^{-1}\!\|\|A(\!\alpha_{i}\!)]_{\mathcal{I}_{i}\mathcal{I}_{i}}^{-1}\|A(\!\alpha_{i+1}\!)\!-\!A(\!\alpha_{i}\!)\|\|\|b(\!\alpha_{i+1}\!)\|\!+\!\|A(\alpha_{i})]_{\mathcal{I}_{i}\mathcal{I}_{i}}^{-1}\|\|b(\alpha_{i+1})\!-\!b(\alpha_{i})\|$
$\textstyle=(\alpha_{i+1}-\alpha_{i})\left(\|A(\alpha_{i+1})]_{\mathcal{I}_{i}\mathcal{I}_{i}}^{-1}\|\|A(\alpha_{i})]_{\mathcal{I}_{i}\mathcal{I}_{i}}^{-1}\|\Delta A\|\|\|b(\alpha_{i+1})\|+\|A(\alpha_{i})]_{\mathcal{I}_{i}\mathcal{I}_{i}}^{-1}\|\|\Delta b\|\right)$
$\textstyle\leq(\alpha_{i+1}-\alpha_{i})\left(\overline{L_{h}}\|A(\alpha_{i+1})]_{\mathcal{I}_{i}\mathcal{I}_{i}}^{-1}\|\|A(\alpha_{i})]_{\mathcal{I}_{i}\mathcal{I}_{i}}^{-1}\|\|b(\alpha_{i+1})\|+L_{f}\|A(\alpha_{i})]_{\mathcal{I}_{i}\mathcal{I}_{i}}^{-1}\|\right)nRr_{2}^{2}$
Further, from Lemma E.1 and Cauchy’s interlacing theorem (cf. [hwang2004cauchy]) we know that for any principal minor of $A(\alpha)$ we have
$\textstyle\|[A(\alpha)^{-1}]_{\mathcal{I}\mathcal{I}}\|\leq\frac{64}{\underline{L}_{h}^{2}},~~\forall\alpha\in[0,1]$
and clearly $\|b(\alpha)\|\leq nL_{f}\overline{L}_{h}+\|K\|H$. And thus we get
$\textstyle\|\lambda(\alpha_{i+1})-\lambda(\alpha_{i})\|\leq(\alpha_{i+1}-\alpha_{i})\left(\frac{4096n\overline{L}_{h}(L_{f}\overline{L}_{h}+\|K\|H)}{\underline{L}_{h}^{4}}+\frac{64L_{f}}{\underline{L}_{h}^{2}}\right)nRr_{2}^{2}$
And thus
$\textstyle\|\lambda-\lambda^{\star}\|=\|\lambda(1)-\lambda(0)\|\leq\left(\frac{4096n\overline{L}_{h}(L_{f}\overline{L}_{h}+\|K\|H)}{\underline{L}_{h}^{4}}+\frac{64L_{f}}{\underline{L}_{h}^{2}}\right)nRr_{2}^{2}.$
@@
$\textstyle\|J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top}(\lambda_{t}^{\star}\!\!-\!\lambda_{t})\|\leq n\overline{L}_{h}^{2}\|\lambda_{t}^{\star}\!-\!\lambda_{t}\|\!\leq\!n^{2}\overline{L}_{h}^{2}R\left(\frac{4096n\overline{L}_{h}(L_{f}\overline{L}_{h}+\|K\|H)}{\underline{L}_{h}^{4}}\!+\!\frac{64L_{f}}{\underline{L}_{h}^{2}}\right)r_{2}^{2}$
Appendix E Bounding ${\lambda_{\min}}(J_{h}(x_{t})\widetilde{J}_{h}(x_{t}))$
Lemma E.1.
In Algorithm 1, we have that given
[…]
then with probability at least $1-\delta$
$\textstyle{\lambda_{\min}}(J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top})\geq\frac{\underline{L}_{h}^{2}}{32}-\overline{L}_{h}Rr_{1}^{2},~~\sigma_{\min}(G_{h})\geq\frac{\underline{L}_{h}^{2}}{32}-\overline{L}_{h}R(r_{1}^{2}+nr_{2}^{2}).$
for all $t=1,2,\dots,T_{G}$
Proof E.2.
From Taylor’s expansion and Assumption 3 we have that
$\textstyle\widetilde{J}_{h}(x_{t})=\frac{n}{T_{B}}\sum_{i=1}^{T_{b}}J_{h}(x_{t})u_{i}u_{i}^{\top}+\epsilon(x_{t}),$
where $\|\epsilon(x_{t})\|\leq Rr_{1}^{2}$. Thus
$\textstyle J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top}=\frac{n}{T_{B}}\sum_{i=1}^{T_{b}}J_{h}(x_{t})u_{i}u_{i}^{\top}J_{h}(x_{t})+\widetilde{\epsilon}(x_{t}),~~~\textup{where }\|\widetilde{\epsilon}(x_{t})\|\leq\overline{L}_{h}Rr_{1}^{2}$
Further, from Lemma E.5 we have that when
[…]
then with probability at least $1-\delta$
$\textstyle{\lambda_{\min}}\left(\frac{n}{T_{B}}\sum_{i=1}^{T_{B}}J_{h}(x_{t})u_{i}u_{i}^{\top}J_{h}(x_{t})\right)\geq\frac{\underline{L}_{h}^{2}}{32},~~\forall~t=1,2,\dots,T_{G}$